I am 47years old and my son is 2years old. How old would I be when I am three times his age?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

I am 47years old and my son is 2years old.
How old would I be when I am three times his age?
----
Solve:
47+x = 3(2+x)
---
47+x = 6+3x
---
2x= 41
x = 20.5
-----
When you are 47+20.5 = 67.5
Your son will be 2+20.5 = 22.5
---
You will then be 3 times your son's age.
===========================================
Cheers,
Stan H.

Answer by nerdybill(7384)

Let x = years from now
then
47+x = 3(2+x)
47+x = 6+3x
47 = 6+2x
41 = 2x
20.5 years = x